OpenStudy (anonymous):

Pls Explain motion in a magnetic field ,when velocity is at an angle with the field .

OpenStudy (anonymous):

??

OpenStudy (anonymous):

Okay, Can you tell me the Magnetic Field Force ? I forgot sorry

OpenStudy (anonymous):

F = q (V x B )

OpenStudy (anonymous):

Hmm Right so do you know cross product ?

OpenStudy (anonymous):

i'm a bit confused about it

OpenStudy (anonymous):

hmm I won't be able to explain through words but I will try finding some videos fro you. btw do you know about mit ocw ?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

what is that?

OpenStudy (anonymous):

oh thanks i'll check it out :)

OpenStudy (anonymous):

if you have any doubts please do post them I have to study magnetic field too so your doubts can be helpful to me. sorry If I sound mean

OpenStudy (anonymous):

no not at all :) .. i'm going through those videos .. it covers most of my portions ...thank you so much

OpenStudy (anonymous):

: ) no problem

OpenStudy (anonymous):

 can you tell me how we get v || =\[v \cos \theta\] and v perpendicular = \[v \sin \theta\]

OpenStudy (anonymous):

are u here ishaan ?

OpenStudy (anonymous):

yea I am ! :D

OpenStudy (anonymous):

can you tell me how we get v || =vcosθ and v perpendicular = vsinθ

OpenStudy (anonymous):

Okay...I will try